OpenStudy (gokuporter):

-1/3x=12 A.-1/4 B.-4 C.-36 D.4

OpenStudy (gokuporter):

B

OpenStudy (anonymous):

Answer is not B

OpenStudy (anonymous):

Multiply by 3 for both sides

OpenStudy (anonymous):

You get 36

OpenStudy (calculusxy):

\[12 \div -\frac{ 1 }{ 3 } = \frac{ 12 }{ 1 } \times \frac{ -3 }{ 1 } = \frac{ -36 }{ 1 } = -36\]

OpenStudy (anonymous):

Then multiply by the negative. Final answer is -36.